OpenStudy (anonymous):

a spinner has 12 congruent sections numbered 1 through 12. if the spinner is spun 1000 times, about how many times would it be expected to land on a number greater than 8?

OpenStudy (anonymous):

A.83 B.167 C.333 D.417

OpenStudy (anonymous):

p = 4/12 N = 1000 Expectation = Np = 1000*4/12 = 333.33 times

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw